Sanchez Company has 38,000 shares of 4% preferred stock of $100 par and 87,000 shares of $50 par common stock. The following amounts were distributed as dividends:

Year 1 $549,000
Year 2 $486,000
Year 3 $526,000

Determine the dividends per share for preferred and common stock for each year. Round the dividends per share to the nearest cent.

Year 1 Year 2 Year 3
Amount distributed $ $ $
Preferred dividend $ $ $
Common dividend $ $ $
Dividends per share:
Preferred stock $ $ $
Common stock $ $ $
Answer #1

ANSWER:

year 1 year 2 year 3
amount distributed $549000 $486000 $526000
preferred dividend

38000 x $4

= $152000

38000 x $4

= $152000

38000 x $4

= $152000

common dividend $397000 $334000 $374000
dividend per share
preferred stock

$100 x 4%

= $4 per share

$100 x 4%

= $4 per share

$100 x 4%

= $4 per share

common stock $397000/87000 = $4.56 per share $334000/87000 = $3.84 per share $374000/87000 = $4.30 per share